OpenStudy (anonymous):

find indicated limit: lim ln cos 4x / 3x^2 x->0 I know ln(u) = u' / u u= cos 4x u' = -4 sinx -4 sinx / cos 4x ?

OpenStudy (anonymous):

Write it using equation tab which is at the bottom left of commenting box?

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{\ln(\cos(4x))}{3x^2}\]

OpenStudy (anonymous):

probably needs l'hopital a couple times

OpenStudy (anonymous):

first time you get \[\frac{-4\sin(4x)}{6x\cos(4x)}\]

OpenStudy (anonymous):

yeah it needs l hospital couple of times

OpenStudy (anonymous):

next one will do it. use the product rule for the denominator

OpenStudy (anonymous):

next one would be \[\frac{ -4\sec^2x }{6}\] now apply limits

OpenStudy (anonymous):

not sure that is right

OpenStudy (anonymous):

the derivative of the numerator is \(16\cos(4x)\) and the derivative of the denominator is \[6(\cos(4x)-4x\sin(4x))\]

OpenStudy (anonymous):

sorry its\[\frac{ -16\sec^2{4x} }{ 6 }\]

OpenStudy (anonymous):

Thnx @satellite73 for pointing it out!

OpenStudy (anonymous):

as \[\frac{ -4\sin(4x) }{ 6xcos(4x) } = \frac{ -4(\sin{4x}/\cos{4x}) }{ 6x }=\frac{ -4\tan4x }{6x}\]

OpenStudy (anonymous):

that is much snappier !!

OpenStudy (anonymous):

yeah @satellite73 i was about to say this is the easy way to work out!

OpenStudy (anonymous):

Thank you guys!